John is the manager at Highway One Electronics. He has been working there for 25 years, and will be retiring at the end of the year. His pension is figured by taking the average of his four highest years of salary. He earned $79,000, $82,000, $83,210, and $84,510 for those four years. His pension plan states that he will get 1.5% of that average for each year he worked. What is his monthly pension amount?

a. $31,691.25 b. $2,568.13 c. $2,640.94 d. $31,290.23


1
Expert's answer
2021-10-04T12:16:46-0400

Average of the highest four years salary

"=\frac{79000+82000+83210+84510}{4}=\$82180"

He worked for 25 years

Total pension "25×\$82180=\$2054500"

Pension of each year="\frac{1.5}{100}×2054500=\$30817.5"

Pension of each month

"=\frac{30817.5}{12}\\=\$2568.125"

Answer

b. $2,568.13